OpenStudy (abbycross167):

How would I put this on a graph? y = -x + 4

OpenStudy (dariusx):

take two values of x for example take x as 0 and you will get y=4. put one dot (0;4) then take x = 2 and you will get y = 2 (2;2) put dots (0;4) and (2;2) on the graph and connect them

OpenStudy (dariusx):

you only need two dots because this is an equation of a slope if it were a parabola equation then you would need a minimum of 3 dots, but for a slope you only need 2

OpenStudy (abbycross167):

okay, umm I forgot to add the directions from my lesson... If I show you them could you help me figure it out? Please? @DariusX

OpenStudy (dariusx):

you mean vectors?

OpenStudy (abbycross167):

Umm no Sir/ma'am... Can I show you the directions?

OpenStudy (dariusx):

sure, let's see what you have there

OpenStudy (abbycross167):

Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. y = -x + 4

OpenStudy (abbycross167):

@DariusX are you able to help me?

OpenStudy (dariusx):

intercepts method is about getting two dots by equating both x and y to 0. first you get one dot by taking x = 0 then from y = -x + 4 you will get: y = 4 that's a dot (0;4) the second dot you will get by taking y = 0 then from y = -x + 4 you will get: -x + 4 = 0 -x = -4 x = 4 that's a dot (4;0) now put those two dots on the graph and connect them

alones (alones):

you need to label it on a graph..i think Darius showed all the steps :)

OpenStudy (abbycross167):

Ok thank you soo much!! But what would the points be that I need to graph? 4 and what other number?

OpenStudy (abbycross167):

@DariusX

OpenStudy (dariusx):

two points (4;0) and (0;4) numbers in brackets represent the dot. left side is x, right side is y looks like this: (x;y) x is horizontal axis y is vertical axis for example: if you have a dot (2;3) then it means x = 2 and y = 3. on the graph the dot will be here